OpenStudy (anonymous):

A train overtakes 2 people who are walking in the same direction as the train is going, at the speed of 2 km/h and 4km/h and the train passes them in 9 & 10 seconds respectively. Find the length and speed of the train.

OpenStudy (anonymous):

Let l be length, s be speed then (s - 5/9)9=l (s - 10/9)10=l

OpenStudy (anonymous):

I did not understand that

OpenStudy (anonymous):

the length of the train, in this case the distance traveled , is the time in overtaking multiplied by the relative velocity

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

him1618 - How do you calculate relative velocity?

OpenStudy (anonymous):

V(A wrt B) = V(A) - V(B) This is a vector equation which can be read as: Velocity vector of A with respect to B equals Velocity vector of A - Velocity vector of B

OpenStudy (anonymous):

Ok - thanks!